OpenStudy (anonymous):

Lori recently rode her bicycle to visit her friend who lives 12 miles away. On her way there, her average speed was 7 miles per hour faster than on her way home. If Lori spent a total of 1 hour bicycling, find the two rates Loris average speed to her friends house: Loris average speed from her friends house:

OpenStudy (anonymous):

i got 21 but idk which number goes where

OpenStudy (calculusxy):

Time = Distance/Rate Time in total = 1 hour To go there: \(\large 1 = \frac{12}{(x+7)}\) To come back: \(\large 1 = \frac{12}{x}\) Add them together to find \(x\) which is the rate. \(\large \frac{12}{(x+7)} + \frac{12}{x} = 1\) Can you solve for x in here?